A 509 g mass oscillates with an amplitude of 11.0 cm on a spring whose spring constant is 21.0 N/m .

Part B

Part complete

Determine the maximum speed.

Part C

Part complete

Determine the total energy.

Express your answer in joules
Answer #1

hetre.

amplitude , A = 11 cm = 0.11 m

mass , m = 509 g = 0.509 kg

spring constant , K = 21 N/m

the angular speed , w = sqrt(K/m)

w = sqrt(21/0.509) = 6.42 rad/s

the maximum speed , Vm = A * w

vm = 0.11 * 6.42 m/s = 0.71 m/s

c)

the total energy , U = 0.5 * K * A^2

U = 0.5 * 21 * 0.11^2 J

U = 0.13 J